answer the following question with regard to the investment theory: a fixed term investment plan will release an amount of $20000 in favor of john in 4 years when he turns 19 and would be ready for enrollment for tertiary studies at a university. naturally, he is curious to know as to what the present value of that amount is. if the discount rate is 7%, kindly assist john to find out what the present value of the $20000 is.
1
Expert's answer
2013-03-29T05:59:40-0400
According to investment theory, if we want to find the present value, we should discount the future value of $20,000 by the
discount rate of 7%.
So, PV = FV/(1 + i)^n, where i - discount rate, n - number of years, PV and FV
- present and future value respectively.
PV = 20,000/1.07^4 = $15,257.9